# Symmetric Linear Bandits with Hidden Symmetry

### v1

@@ Abstract
High-dimensional linear bandits with low-dimensional structure have received considerable attention in recent studies due to their practical significance. The most common structure in the literature is sparsity. However, it may not be available in practice. Symmetry, where the reward is invariant under certain groups of transformations on the set of arms, is another important inductive bias in the high-dimensional case that covers many standard structures, including sparsity. In this work, we study high-dimensional symmetric linear bandits where the symmetry is hidden from the learner, and the correct symmetry needs to be learned in an online setting. We examine the structure of a collection of hidden symmetry and provide a method based on model selection within the collection of low-dimensional subspaces. Our algorithm achieves a regret bound of $O(d_{0}^{1/3}T^{2/3}\log(d))$, where $d$ is the ambient dimension which is potentially very large, and $d_{0}$ is the dimension of the true low-dimensional subspace such that $d_{0}\ll d$. With an extra assumption on well-separated models, we can further improve the regret to $O(d_{0}\sqrt{T\log(d)})$.
@@ 1 Introduction
Stochastic bandit is a sequential decision-making problem in which a player, who aims to maximize his reward, selects an action at each step and receives a stochastic reward, drawn from an initially unknown distribution of the selected arm, in response. Linear stochastic bandit (LSB) [2] is an important variant in which the expected value of the reward is a linear function of the action. It is one of the most studied bandit variants and has many practical applications [24].
Actions in LSB are specified as feature vectors in $\mathbb{R}^{d}$ for very large feature dimension $d$, with performance i.e. the resulting regret scaling with $d$. Many works have addressed this curse of dimensionality by leveraging different low-dimensional structures as inductive biases for the learner. For example, sparsity, which assumes that the reward is a sparse linear function, has been used extensively in LSB to design bandit algorithms with better performance [3, 32, 21]. However, when the reward function lacks the structure fo sparsity (which may occur in many real-world situations), a question arises: Are there different structures in the features of LSB that we can exploit to overcome the curse of dimensionality and design bandit algorithms with better performance?
@@
[…]
Against this background, we ask the question of whether learner can leverage symmetry to enable effective exploration, and break the curse of dimensionality without knowledge of the symmetry structure? Moreover, in the presence of symmetry, when can we design learning algorithms with optimal regret bounds? Towards answering this question, we investigate the setting of symmetric linear stochastic bandit in $\mathbb{R}^{d}$, where $d$ is potentially very large, and the expected reward function is invariant with respect to the actions of a hidden group $\mathcal{G}$ of coordinate permutations. Our contributions are summarised as follows:
@@
    Given this, we establish a cardinality condition on the class of symmetric linear bandits with hidden $\mathcal{G}$, in which the learner can learn the true symmetry structure and overcome the curse of dimensionality. Notably, this class includes sparsity as a special case, and therefore inherits all the computational and statistical complexities of sparsity.
@@
    We cast our problem of learning with hidden subgroup $\mathcal{G}$ into model selection with collection of low-dimensional subspaces [22, 31]. To address the polynomial scaling of regret bounds with respect to the number of models and arms in previous works, we depart from model aggregation, which is typically used in LSB model selection, and introduce a new framework inspired by Gaussian model selection [7]  [^1] We note that “Gaussian model selection” is a technique in statistics, similar to model aggregation (see [19]’s chapter 2 and 4), which should not be confused with “model selection” in the bandit literature. and compressed sensing [8]. Based on this framework, we introduce a new algorithm, called EMC (for Explore Models then Commit). Under the assumption that the set of arm is exploratory, we prove that the regret bound of the EMC algorithm is $O(d_{0}^{1/3}T^{2/3}\log(d))$, and $O(d_{0}\sqrt{T}\log(d))$ with an additional assumption on well-separated partitions, where $d_{0}\ll d$ is the dimension of the low-dimensional subspace associated with group $\mathcal{G}$.
@@ 1.1 Related Work
We now briefly outline related work and compare them with our results. We refer the reader to Appendix D.2 for a more in-depth literature review.
@@
Sparse linear bandits. As we will explain in Section 4.2, sparsity is equivalent to a subset symmetry structures, and thus, can be seen as a special case of our setting. As such, we first review the literature of sparsity. Sparse linear bandits were first investigated in [3], where the authors achieve a regret of $\tilde{O}(\sqrt{dsT})$, with $\tilde{O}$ disregarding the logarithmic factor, and $s$ representing the sparsity level, and $T$ is the time horizon. This matches the regret lower bound for sparse bandits, which is $\Omega(\sqrt{dsT})$ [24]. More recently, the contextual version of linear bandits has gained popularity, where additional assumptions are made regarding the context distribution and set of arms [23, 32, 25, 11, 21] to avoid polynomial dependence on $d$. Notably, with the assumption on exploratory set of arms, [21] propose an Explore then Commit style strategy that achieves $\tilde{O}(s^{\frac{2}{3}}T^{\frac{2}{3}})$, nearly matching the regret lower bound $\Omega(s^{\frac{1}{3}}T^{\frac{2}{3}})$ in the data-poor regime. As sparsity is equivalent to a subclass of hidden symmetry, all the lower bounds for sparse problems apply to our setting of learning with hidden symmetry.
@@
Model selection. Our problem is also closely related to the problem of model selection in linear bandits, as the learner can collect potential candidates for the hidden symmetry model. Particularly, in model selection, there is a collection of $M$ features, and different linear bandits running with each of these features serve as base algorithms. By exploiting the fact that the data can be shared across all the base algorithms, the dependence of regret in terms of the number of features can be reduced to $\log(M)$. In particular, [22] propose a method that concatenates all $M$ features of dimension $d$ into one feature of dimension $Md$, and uses the Lasso estimation as a aggregation of models. Their algorithm achieves a regret bound of $O(T^{\frac{3}{4}}\sqrt{\log(M)})$ under the assumption that the Euclidean norm of the concatenated feature is bounded by a constant. However, in our case, the Euclidean norm of the concatenated feature vector can be as large as $\sqrt{M}$, which leads to a $\sqrt{M}$ multiplicative factor in the regret bound. Besides, [31] uses the online aggregation oracle approach, and able to obtain regret of $O(\sqrt{KdT\log(M)})$, where $K$ is the number of arms. In contrast, we use algorithmic mechanisms that are different to aggregation of models. In particular, we explicitly exploit the structure of the model class as a collection of subspaces and invoke results from Gaussian model selection [19, 7] and dimension reduction on the union of subspaces [8]. With this technique, we are able to achieve $O(T^{\frac{2}{3}}\log(M))$, which is rate-optimal in the data-poor regime, has logarithmic dependence on $M$ without strong assumptions on the norm of concatenated features, and is independent of the number of arms $K$. We refer the reader to Section 4.2 for a more detailed explanation.
@@
Symmetry in online learning. The notion of symmetry in Markov Decision Making dates back to works such as [20, 36]. Generally, the reward function and probability transition are preserved under an action of a group on the state-action space. Exploiting known symmetry has been shown to help achieve better performance empirically [39, 38] or tighter regret bounds theoretically [37]. However, all these works requires knowledge of symmetry group, while our setting consider hidden symmetry group which may be considerably harder. Hidden symmetry on the context or state space has been studied by few authors, with the term context-lumpable bandits [26], meaning that the set of contexts can be partitioned into classes of similar contexts. It is important to note that the symmetry group acts differently on the context space and the action space. As we shall explain in detail in Section 3, while one can achieve a reduction in terms of regret in the case of hidden symmetry acting on context spaces [26], this is not the case when the symmetry group acts on the action space. The work closest to ours is [35], where the authors consider the setting of a $K$-armed bandit, where the set of arms can be partitioned into groups with similar mean rewards, such that each group has at least $q>2$ arms. With the constrained partition, the instance-dependent regret bounds are shown asymptotically to be of order $O\left(\tfrac{K}{q}\log T\right)$. Comparing to [35], we study the setting of stochastic linear bandits with similar arms, in which the (hidden) symmetry and linearity structure may intertwine, making the problem more sophisticated. We also impose different constraints on the way one partitions the set of arms, which is more natural in the setting of linear bandits with infinite arms.
@@ 2 Problem Setting
ϕ^::^italic-$\hat{\phi}:$ $\mathcal{S}_{d}\times\mathbb{R}^{d}$ $\rightarrow\mathbb{R}^{d}$ (1)
@@
In other words, a group element $g$ acts on an arm $x\in\mathbb{R}^{d}$ by permuting the coordinates of $x$. In the setting of linear bandit, the permutation group action also acts on the set of parameters via coordinate permutation as
$\left(g,(\theta_{i})_{i\in[d]}\right)$ $\mapsto(\theta_{g(i)})_{i\in[d]}$
For brevity, we simply denote $g\cdot\theta$ and $g\cdot x$ as $\phi(g,\theta)$ and $\hat{\phi}(g,x)$, respectively. Denote by $A_{g}$ the permutation matrix corresponding to $g$. We write $\mathcal{G}\leq\mathcal{S}_{d}$ to denote that $\mathcal{G}$ is a subgroup of $\mathcal{S}_{d}$. Given any point $\theta\in\mathbb{R}^{d}$, we write $\mathcal{G}\cdot\theta=\left\{g\cdot\theta\mid g\in\mathcal{G}\right\}$ to denote the orbit of $\theta$ under $\mathcal{G}$. It is well known that the orbit induced by the induced action of a subgroup $\mathcal{G}\leq\mathcal{S}_{d}$ corresponds to a set partition of $[d]$. We denote this partition as $\pi_{\mathcal{G}}$.
@@ 3.1 Fixed Point Subspace and Partition
Given $d\in\mathbb{N}^{+}$, we denote $\mathcal{P}_{d}$ be the set of all partitions of $[d]$. Let $\mathcal{P}_{d,k}$ as the set of all partitions of $[d]$ with exactly $k$ classes, and $\mathcal{P}_{d,\leq k}$ be the set of all partitions of $[d]$ with at most $k$ classes. The number of set partitions with $k$ classes $|\mathcal{P}_{d,k}|$ is known as the Stirling number of the second kind, and $|\mathcal{P}_{d}|$ is known as Bell number.
@@ 3.2 Impossibility Result
By Proposition 2, we have that the number of subspaces of dimension $k$ in $\mathcal{F}_{\mathcal{S}_{d}}$ is exactly the number of set partitions with $k$-classes. Suppose that the learner knows that the orbit under action of $\mathcal{G}$ partitions the index of $\theta_{\star}$ into 2 equivalent classes that is, $\mathrm{dim}(\mathrm{Fix}_{\mathcal{G}})=2$. The learner cannot get any reduction in term of regret.
@@
Assume that the action set is the unit ball $\mathcal{X}=\{x\in\mathbb{R}^{d}\mid\left\|x\right\|_{2}\leq 1\}$, and $f$ is invariant w.r.t. action of subgroup $\mathcal{G}\leq\mathcal{S}_{d}$, such that $\mathrm{dim}(\mathrm{Fix}_{\mathcal{G}})=2$. Then, for any $T\geq d/2$, regret of any bandit algorithm is lower bounded by $\mathbf{R}_{T}=\Omega(d\sqrt{T})$.
@@ 4 The Case of Hidden Subgroups with Subexponential Size
As indicated by Proposition 3, there is no improvement in terms of regret, despite the learner having access to a collection of extremely low-dimensional fixed point subspaces. Therefore, we assume that the learner can access only a reasonably small subset of the collection of low-dimensional fixed point subspaces. Let $d_{0}$ be the upper bound for the dimension of fixed point subspaces; that is, we know that the orbit of $\mathcal{G}$ partitions $[d]$ into at most $d_{0}$ classes. Now, let us assume that the learner knows that $\mathcal{G}$ does not partition $[d]$ freely, but must satisfy certain constraints, that is, $\pi_{\mathcal{G}}\in\mathcal{Q}_{d,\leq d_{0}}\subset\mathcal{P}_{d,\leq d_{0}}$. Here, $\mathcal{Q}_{d,\leq d_{0}}$ is a small collection of partitions with at most $d_{0}$ classes, which encodes the constraints on the way $\mathcal{G}$ partitions $[d]$. We introduce an assumption regarding the cardinality of $\mathcal{Q}_{d,\leq d_{0}}$, which are formally stated in Section 4.2. Using the Proposition 1, we can define the collection of fixed point subspaces associated with the collection of partition $\mathcal{Q}_{d,\leq d_{0}}$ via the bijection $\mathbf{H}$ as

### v2

@@ Abstract
High-dimensional linear bandits with low-dimensional structure have received considerable attention in recent studies due to their practical significance. The most common structure in the literature is sparsity. However, it may not be available in practice. Symmetry, where the reward is invariant under certain groups of transformations on the set of arms, is another important inductive bias in the high-dimensional case that covers many standard structures, including sparsity. In this work, we study high-dimensional symmetric linear bandits where the symmetry is hidden from the learner, and the correct symmetry needs to be learned in an online setting. We examine the structure of a collection of hidden symmetry and provide a method based on model selection within the collection of low-dimensional subspaces. Our algorithm achieves a regret bound of $O(d_{0}^{2/3}T^{2/3}\log(d))$, where $d$ is the ambient dimension which is potentially very large, and $d_{0}$ is the dimension of the true low-dimensional subspace such that $d_{0}\ll d$. With an extra assumption on well-separated models, we can further improve the regret to $O(d_{0}\sqrt{T\log(d)})$.
@@ 1 Introduction
Stochastic bandit is a sequential decision-making problem in which a player, who aims to maximize her reward, selects an action at each step and receives a stochastic reward, drawn from an initially unknown distribution of the selected arm, in response. Linear stochastic bandit (LSB) [1] is an important variant in which the expected value of the reward is a linear function of the action. It is one of the most studied bandit variants and has many practical applications [27].
Actions in LSB are specified as feature vectors in $\mathbb{R}^{d}$ for very large feature dimension $d$, with performance i.e. the resulting regret scaling with $d$. Many works have addressed this curse of dimensionality by leveraging different low-dimensional structures as inductive biases for the learner. For example, sparsity, which assumes that the reward is a sparse linear function, has been used extensively in LSB to design bandit algorithms with better performance [2, 36, 23]. However, when the reward function lacks the structure for sparsity (which may occur in many real-world situations), a question arises: Are there different structures in the features of LSB that we can exploit to overcome the curse of dimensionality and design bandit algorithms with better performance?
@@
In the machine learning literature, especially in supervised learning, most studies on symmetry assume prior knowledge of symmetry structures of the tasks under consideration [31, 16]. However, in numerous practical scenarios, the learner can only access to partial knowledge of the symmetry, necessitating the incorporation of symmetry learning mechanisms into the algorithms to achieve better performance. Examples of hidden symmetry can be found in multi-agent learning with cooperative behavior. As a motivating example, consider a company undertaking a large project that consists of several subtasks. The company must hire subcontractors with the goal of maximizing project quality while staying within budget constraints. Symmetry may arise in this situation when coalitions form among subcontractors, where members of a coalition work together to complete their allocated tasks using shared resources. In particular, the allocation of tasks within a coalition can be swapped without affecting overall team performance, inducing symmetry (i.e., performance remains invariant under permutation) in the task assignments. Coalitions among subcontractors often arise since sharing labor and resources reduces operational costs, making their work more efficient and cost-effective. However, these coalitions are typically hidden from the hiring company. One reason is that if the hiring company were aware of these collaborations, they could use this information to negotiate lower prices, knowing that the subcontractors are benefiting from shared resources. Another reason is that coalitions may raise concerns about collusion. In particular, in a competitive market, such as when subcontractors are hired through a bidding platform, coalition members can collaborate to manipulate the bidding process, which is considered unfair and could undermine the integrity of the bidding. For more practical examples of hidden symmetry in multi-agent reinforcement learning [34] and robotics [3], we refer the reader to Appendix D.2. Motivated by these examples, we believe that hidden symmetry is much more relevant in the context of sequential decision-making because the environment and its symmetry structure may not be readily available to the learner, as opposed to supervised learning and offline settings where data are provided during the training phase. As the learner has the power to freely collect data, it is expected that they will learn the hidden symmetry structures as they explore the environment.
Against this background, we ask the question of whether learner can leverage symmetry to enable effective exploration, and break the curse of dimensionality without a prior knowledge of the symmetry structure? Moreover, in the presence of symmetry, when can we design learning algorithms with optimal regret bounds? Towards answering this question, we investigate the setting of symmetric linear stochastic bandit in $\mathbb{R}^{d}$, where $d$ is potentially very large, and the expected reward function is invariant with respect to the actions of a hidden group $\mathcal{G}$ of coordinate permutations. Our contributions are summarised as follows:
@@
    Given this, we establish a cardinality condition on the class of symmetric linear bandits with hidden $\mathcal{G}$, in which the learner can learn the true symmetry structure and overcome the curse of dimensionality. Notably, this class includes a sparsity as a special case, and therefore inherits all the computational and statistical complexities of sparsity. Apart from sparsity, this class includes many other practically relevant classes, such as partitions that respect underlying hierarchies, non-crossing partitions, and non-nesting partitions (see Subsection 4.2.1 and Appendix D.2).
@@
    We cast our problem of learning with hidden subgroup $\mathcal{G}$ into model selection with collection of low-dimensional subspaces [25, 35]. To address the polynomial scaling of regret bounds with respect to the number of models and arms in previous works, we depart from model aggregation, which is typically used in LSB model selection, and introduce a new framework inspired by Gaussian model selection [7]  [^1] We note that “Gaussian model selection” is a technique in statistics, similar to model aggregation (see [21]’s chapter 2 and 4), which should not be confused with “model selection” in the bandit literature. and compressed sensing [8]. Based on this framework, we introduce a new algorithm, called EMC (for Explore Models then Commit). Under the assumption that the set of arm is exploratory, we prove that the regret bound of the EMC algorithm is $O(d_{0}^{2/3}T^{2/3}\log(d))$, and $O(d_{0}\sqrt{T}\log(d))$ with an additional assumption on well-separated partitions, where $d_{0}\ll d$ is the dimension of the low-dimensional subspace associated with group $\mathcal{G}$.
@@ 1.1 Related Work
We now briefly outline related work and compare them with our results. We refer the reader to Appendix F for a more in-depth literature review.
@@
Sparse linear bandits. As we will explain in Section 4.2, sparsity is equivalent to a subset symmetry structures, and thus, can be seen as a special case of our setting. As such, we first review the literature of sparsity. Sparse linear bandits were first investigated in [2], where the authors achieve a regret of $\tilde{O}(\sqrt{dsT})$, with $\tilde{O}$ disregarding the logarithmic factor, and $s$ representing the sparsity level, and $T$ is the time horizon. This matches the regret lower bound for sparse bandits, which is $\Omega(\sqrt{dsT})$ [27]. More recently, the contextual version of linear bandits has gained popularity, where additional assumptions are made regarding the context distribution and set of arms [26, 36, 28, 11, 23] to avoid polynomial dependence on $d$. Notably, with the assumption on exploratory set of arms, [23] propose an Explore then Commit style strategy that achieves $\tilde{O}(s^{\frac{2}{3}}T^{\frac{2}{3}})$, nearly matching the regret lower bound $\Omega(s^{\frac{2}{3}}T^{\frac{2}{3}})$ [24] in the data-poor regime. As sparsity is equivalent to a subclass of hidden symmetry, all the lower bounds for sparse problems apply to our setting of learning with hidden symmetry.
@@
Model selection. Our problem is also closely related to the problem of model selection in linear bandits, as the learner can collect potential candidates for the hidden symmetry model. Particularly, in model selection, there is a collection of $M$ features, and different linear bandits running with each of these features serve as base algorithms. By exploiting the fact that the data can be shared across all the base algorithms, the dependence of regret in terms of the number of features can be reduced to $\log(M)$. In particular, [25] propose a method that concatenates all $M$ features of dimension $d$ into one feature of dimension $Md$, and uses the Lasso estimation as a aggregation of models. Their algorithm achieves a regret bound of $O(T^{\frac{3}{4}}\sqrt{\log(M)})$ under the assumption that the Euclidean norm of the concatenated feature is bounded by a constant. However, in our case, the Euclidean norm of the concatenated feature vector can be as large as $\sqrt{M}$, which leads to a $\sqrt{M}$ multiplicative factor in the regret bound. Besides, [35] uses the online aggregation oracle approach, and is able to obtain regret of $O(\sqrt{KdT\log(M)})$, where $K$ is the number of arms. In contrast, we use algorithmic mechanisms that are different from aggregation of models. In particular, we explicitly exploit the structure of the model class as a collection of subspaces and invoke results from Gaussian model selection [21, 7] and dimension reduction on the union of subspaces [8]. With this technique, we are able to achieve $O(T^{\frac{2}{3}}\log(M))$, which is rate-optimal in the data-poor regime, has logarithmic dependence on $M$ without strong assumptions on the norm of concatenated features, and is independent of the number of arms $K$. We refer the reader to Section 4.2 for a more detailed explanation.
@@
Symmetry in online learning. The notion of symmetry in Markov decision process dates back to works such as [22, 40]. Generally, the reward function and probability transition are preserved under an action of a group on the state-action space. Exploiting known symmetry has been shown to help achieve better performance empirically [43, 42] or tighter regret bounds theoretically [41]. However, all these works requires knowledge of symmetry group, while our setting consider hidden symmetry group which may be considerably harder. Hidden symmetry on the context or state space has been studied by few authors, with the term context-lumpable bandits [29], meaning that the set of contexts can be partitioned into classes of similar contexts. It is important to note that the symmetry group acts differently on the context space and the action space. As we shall explain in detail in Section 3, while one can achieve a reduction in terms of regret in the case of hidden symmetry acting on context spaces [29], this is not the case when the symmetry group acts on the action space. The work closest to ours is [39], where the authors consider the setting of a $K$-armed bandit, where the set of arms can be partitioned into groups with similar mean rewards, such that each group has at least $q>2$ arms. With the constrained partition, the instance-dependent regret bounds are shown asymptotically to be of order $O\left(\tfrac{K}{q}\log T\right)$. Comparing to [39], we study the setting of stochastic linear bandits with similar arms, in which the (hidden) symmetry and linearity structure may intertwine, making the problem more sophisticated. We also impose different constraints on the way one partitions the set of arms, which is more natural in the setting of linear bandits with infinite arms.
@@ 2 Problem Setting
In other words, a group element $g$ acts on an arm $x\in\mathbb{R}^{d}$ by permuting the coordinates of $x$. In the setting of linear bandit, the permutation group action also acts on the set of parameters via coordinate permutation. For brevity, we simply denote $g\cdot\theta$ and $g\cdot x$ as $\phi(g,\theta)$ and $\phi(g,x)$, respectively. Denote by $A_{g}$ the permutation matrix corresponding to $g$. We write $\mathcal{G}\leq\mathcal{S}_{d}$ to denote that $\mathcal{G}$ is a subgroup of $\mathcal{S}_{d}$. Given any point $\theta\in\mathbb{R}^{d}$, we write $\mathcal{G}\cdot\theta=\left\{g\cdot\theta\mid g\in\mathcal{G}\right\}$ to denote the orbit of $\theta$ under $\mathcal{G}$. It is well known that the orbit induced by the induced action of a subgroup $\mathcal{G}\leq\mathcal{S}_{d}$ corresponds to a set partition of $[d]$. We denote this partition as $\pi_{\mathcal{G}}$.
@@ 3.1 Fixed Point Subspace and Partition
Given $d\in\mathbb{N}^{+}$, we denote $\mathcal{P}_{d}$ as the set of all partitions of $[d]$. Let $\mathcal{P}_{d,k}$ as the set of all partitions of $[d]$ with exactly $k$ classes, and $\mathcal{P}_{d,\leq k}$ be the set of all partitions of $[d]$ with at most $k$ classes. The number of set partitions with $k$ classes $|\mathcal{P}_{d,k}|$ is known as the Stirling number of the second kind, and $|\mathcal{P}_{d}|$ is known as Bell number.
@@ 3.2 Impossibility Result
By Proposition 2, we have that the number of subspaces of dimension $k$ in $\mathcal{F}_{\mathcal{S}_{d}}$ is exactly the number of set partitions with $k$-classes. Suppose that the learner knows that the orbit under action of $\mathcal{G}$ partitions the index of $\theta_{\star}$ into 2 equivalent classes that is, $\mathrm{dim}(\mathrm{Fix}_{\mathcal{G}})=2$. The learner cannot get any reduction in terms of regret.
@@
Assume that the action set is the unit cube $\mathcal{X}=\{x\in\mathbb{R}^{d}\mid\|x\|_{\infty}\leq 1\}$, and $f$ is invariant w.r.t. action of subgroup $\mathcal{G}\leq\mathcal{S}_{d}$, such that $\mathrm{dim}(\mathrm{Fix}_{\mathcal{G}})=2$. Then, the regret of any bandit algorithm is lower bounded by $\mathbf{R}_{T}=\Omega(d\sqrt{T})$.
@@ 4 The Case of Hidden Subgroups with Subexponential Size
As indicated by Proposition 3, there is no improvement in terms of regret, despite the learner having access to a collection of extremely low-dimensional fixed point subspaces. Therefore, we assume that the learner can access only a reasonably small subset of the collection of low-dimensional fixed point subspaces. Let $d_{0}$ be the upper bound for the dimension of fixed point subspaces; that is, we know that the orbit of $\mathcal{G}$ partitions $[d]$ into at most $d_{0}$ classes. Now, let us assume that the learner knows that $\mathcal{G}$ does not partition $[d]$ freely, but must satisfy certain constraints, that is, $\pi_{\mathcal{G}}\in\mathcal{Q}_{d,\leq d_{0}}\subset\mathcal{P}_{d,\leq d_{0}}$. Here, $\mathcal{Q}_{d,\leq d_{0}}$ is a small collection of partitions with at most $d_{0}$ classes, which encodes the constraints on the way $\mathcal{G}$ partitions $[d]$. We introduce an assumption regarding the cardinality of $\mathcal{Q}_{d,\leq d_{0}}$, which is formally stated in Section 4.2. Using the Proposition 1, we can define the collection of fixed point subspaces associated with the collection of partition $\mathcal{Q}_{d,\leq d_{0}}$ via the bijection $\mathbf{H}$ as
@@
In addition, let us define the extension of the collection $\mathcal{M}$ as ¯:=$\,\overline{\!{\mathcal{M}}}:=\left\{\mathrm{conv}\left(m\cup m^{\prime}\right)\mid\;m,\;m^{\prime}\in\mathcal{M}\right\},$ where $\mathrm{conv}(S)$ is the convex hull of the set $S\subset\mathbb{R}^{n}$. We have that $\,\overline{\!{\mathcal{M}}}$ is a collection of subspaces, that is, $\mathrm{conv}\left(m\cup m^{\prime}\right)$ is indeed a subspace [8]. Denote $\,\overline{\!{M}}:=|\,\overline{\!{\mathcal{M}}}|$, then we have $\,\overline{\!{M}}=(M^{2}-M)/2$. Moreover, if dimension of subspace in $\mathcal{M}$ is at most $d_{0}$, then the dimension of subspace in $\,\overline{\!{\mathcal{M}}}$ is at most $2d_{0}$.